OpenStudy (anonymous):

Let u = (1 , 1 , 0) , v = (0 , 1 , 0) and w = (-4 , -26 , -24) . Evaluate the expression || 6u - 4v + w || =

OpenStudy (amistre64):

fist scale the vetors, then add and subtract as needed ... then the length comes in

OpenStudy (amistre64):

*first ...

OpenStudy (amistre64):

6u = 6( 1, 1, 0 ) -4v = -4( 0, 1, 0 ) +w = (-4,-26,-24) ( 6, 6, 0 ) ( 0, -4, 0 ) (-4,-26,-24) ------------- (2, -24, -24) , then lengthify it

OpenStudy (anonymous):

thanks. the answer must bu in the form of ll6u-4v+wll

OpenStudy (anonymous):

*but

OpenStudy (anonymous):

ok. i got it.